Let A = PDP-1 and P and D as shown below. Compute A+ A4-100 00 (Simplify your answers.)
Answer #1

det CP) = 1x3 - 2x2 = 3-4 =-1 co-factors. G1 = 3, C9,5-2, 42=-2 C22=1 part de ces cenas [37] ༥ A = PD po! = (PDP-7 2 = PDPI P